OpenStudy (anonymous):

Solve for x: 7 over 8 minus 1 over x equals 3 over 4 x = 1 x = 2 x = 4 x = 8

OpenStudy (anonymous):

(7/8) - (1/x) = (3/4) first set the equation equal to zero by subtracting 3/4 from both sides (7/8)-(1/x)-(3/4)=0 next find the least common denominator (LCD), in this case, the LCD is 8x multiply each individual term by the LCD (8x) (7/8)*8x = 7x, the 8 is on both the top and the bottom, so it cancels out, the x is multiplied by the 7 because they are both on top (1/x)*8x = 8, the x is on both the top and the bottom, so it cancels out, the 8 is multiplied by the 1 because they are both on top (3/4)*8x = 6x, 8 divided by four equals two, because the greater number is on the top, the quotient is on the top, the x does not cancel out, so it is multiplied by the 2 to get 2x, which is then multiplied by the 3 to get 6x. You could also look at it like (3/4)*(8x/1), which equals 24x/4, or, again, 6x. Now you have: 7x-8-6x = 0 combine like terms (the 7x and -6x) to get: x-8 = 0 add eight to both sides x = 8